OpenStudy (anonymous):

Determine the sum of the given probabilities. Are the events complementary? P(A) = ; P(B) = A. 5/14 The events are complementary. B. 5/14 The events are not complementary. C. 5/7 The events are complementary. D. 5/7 The events are not complementary.

OpenStudy (anonymous):

@TeenWolfGirl

OpenStudy (anonymous):

@JTijerina

OpenStudy (anonymous):

sorry i can't help:(

OpenStudy (displayerror):

Are there supposed to be pictures or some expression after P(A) and P(B)? You haven't provided us with the given probabilities.

OpenStudy (anonymous):

sorry one min

OpenStudy (anonymous):

P(A)= 3/7; P(B)= 2/7

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

Oh, sorry, I am horrible at probabilities.

OpenStudy (anonymous):

@aaliyahfairgood

OpenStudy (anonymous):

okay :(

OpenStudy (anonymous):

can anyone help medal and fan!!!

OpenStudy (displayerror):

I think it is choice D. If we add the probabilities, we get: \[\text{P(A) + P(B)} = \frac{3}{7} + \frac{2}{7} = \frac{5}{7} \neq 1\] If two events are complementary, then that means the sum of their probabilities should equal 1 (we say that the complement of event A is 1 - P(B) or that the complement of event B is 1 - P(A)). Now from adding the two probabilities, we see that the sum doesn't equal to 1, thus we can conclude that the two events are not complementary.

OpenStudy (anonymous):

okay, thank you!